OpenStudy (anonymous):

expression that is equal to ln(10 x) - ln(2 x)?

OpenStudy (anonymous):

is it ln(10 x)/ln(2 x) ?

OpenStudy (anonymous):

= ln(10) + ln(x) - ln(2) - ln(x) =ln(5) (because ln(ab) = ln(a) + ln(b)

OpenStudy (anonymous):

what if it was ln(9 x^2)/(ln (3 x))?

OpenStudy (anonymous):

qoutient rule? right

OpenStudy (anonymous):

= ln((3x)^2) / ln(3x) =2 * ln(3x) / ln(3x) =2 because ln(a^b) = b*ln(a)

OpenStudy (anonymous):

ok thanx

OpenStudy (anonymous):

you're welcome